A consumer advocate claims that 80 percent of cable television subscribers are not satisfied with their cable service. In an attempt to justify this claim, a randomly selected sample of cable subscribers will be polled on this issue. Suppose that the advocate’s claim is true, and suppose that a random sample of five cable subscribers is selected. Assuming independence, use an appropriate formula to compute the probability that four or more subscribers in the sample are not satisfied with their service.

Minitab instructions: Go to Calc > select Probability Distributions > select Binomial > select Cumulative probability > Number of Trials insert 5 > Event Probability insert “.8” > Input Constant insert “3.” Click OK.

Paste your Minitab results here and then show your work to calculate P(x ≥ 4) = 1 – P(x ≤ 3)
Answer #1

from minitab:

output:

Cumulative Distribution Function

Binomial with n = 5 and p = 0.8

x P( X ≤ x )
3 0.26272

therefore

P(x ≥ 4) = 1 – P(x ≤ 3) =1-0.26272 =0.73728